OpenStudy (anonymous):

please help me solve the quadratics using factoring 2x^2 + 4x + 6 = 0 x^2 + 6x + 9 = 9

jhonyy9 (jhonyy9):

2x2 +4x +6 =(x-(-1-isqrt2))(x-(-1+isqrt2)) x2 +6x +9 =(x+3)(x+3)